What is the slope of 4x-3y=36.
Answer by JenniferTutors(83)

To find the slope of 4x-3y=36, put it into this form: y=mx+b
m=slope
b=y-intercept
4x-3y=36
Get y by itself by first subtracting the 4x from both sides of the equation
4x-3y=36
-4x -4x
-3y=-4x+36
Next, to get y completely on its own, divide -3 to both sides of the equation
-3y/-3 = -4x/-3 + 36/-3
y = 4x/3 - 12
slope (m) = 4/3